A jetliner flying with the wind traveled 3600 km in 6 hours. The return trip, against the wind, took 8 hours. Find the rate at which the jet flew in still air.

Answer by Paul(988)

Let the speed of the plane be x
Let the speed of the wind be y

6(x+y)=3600 (solve for y)
6x+6y=3600
6y=3600-6x
y=600-x (subsitution)

8(x-y)=3600
8x-8y=3600 (apply subsitution)
8x-8(600-x)=3600
8x+8x-4800=3600
16x=8400
x=525

y=600-525
y=75

Hence, the speed of the plane in still air is 525mph and the speed of the wind is 75mph.
PAUL.